EA011C

PI: Alef
Title: Cyg A

Session: Feb 1997 (session 1)
Frequency: 18 cm
Mode: A/PI
Obs time (UT): EU 01:30(02/03)-13:30(02/03)
Stations: EbJbMcNtOnTrSmShUrWb
Processor: Bonn
EU Mk3 Tapes/St.: 3

Comment


To give feedback on this experiment, click here

Experiment Feedback

Onsala (Antonis Polatidis at Sun Mar 2 08:42:07 EST 1997) Success
High winds, close to the telescope stowing limit. Tracking was fine though. No known problems. Rain for the second part of the experiment.

Medicina (Daniele Dallacasa at Sun Mar 2 08:54:11 EST 1997) Minor failures
Various time-variable RFI, particularly in VC06 and VC08. They were occasionally saturated. Attemped various attenuators in order to obtain a decent signal. The sky, initially cludy, was clear for most of the experiment.

Effelsberg (dgraham at Tue Mar 4 03:05:08 EST 1997) Minor failures
may be problem with VC11. Tsys measurements on Cyg A no good.

Westerbork (Tony Foley at Tue Mar 4 07:05:25 EST 1997) Success
Looks OK

Shanghai (Huang Xinyong at Wed Mar 5 05:17:43 EST 1997) Success
Clock offset = 0.61-0.63 msec. No problems, weather cloudy.

Jodrell (Alastair Gunn at Tue Mar 18 15:11:46 MET 1997) Minor failures
Telescope parked until 06:40 due to high winds. Parked again at 10:39 due to winds. Very bad interference.

Calibration (Fredrik T Rantakyro at Thu May 8 11:38:53 METDST 1997) Success
The ANTAB file for the experiment has been placed on VLBEER in the Feb97 directory with the name ea11c.antab. This file contains the preliminary calibration results and may be updated when I receive updated information from the telescopes and the results from the datareduction of the calibration experiments.

MPIfR (Walter Alef at Wed Jan 7 12:10:59 MET 1998) Failure
Correlation cancelled due to failures at too many stations.